a company has agreed to pay $2.1 million in 5 years to settle a lawsuit. how much must in invest in an account paying 9% annual intetest compounded monthly to have that amount when is due?
what is the interest rate per period ?
Answer #1

Answer:

Due Amount in 5 years F=2.1 million

annual rate =9%

Period rate =Monthly rate r=9%/12=0.75%

Number of months N=5*12=60

Amount deposited = F/(1+r)^N=2.1/(1+0.75%)^60=$1.3412 Millions